Figure 1: The alt-az coordinate system

This is a diagram which explains the altitudeazimuth, or “alt-az”, coordinate system. It is a purely local system: the coordinates determined by an observer in New York will not agree with those determined simultaneously by an observer in Texas.

Notice that, in the final typeset document, the diagram doesn’t always end up in the same place relative to the surrounding text as it does in the input file.